Mariota Corp. just paid a dividend of $3.45 per share on its stock. The dividend growth rate is expected to be 4.05 forever and investors require a return of 11.9 percent on this stock. What will the stock price be in 11 years?
Answer #1

Dividend per Share =3.45
Growth =4.05%
Stock Price will be in 11 years =Dividend in year 12/(Required Rate-growth) =3.45*(1+4.05%)^12/(11.9%-4.05%) =70.77